# Comparison results for Markov tree distributions

### v1

@@ 1.1 Main result
    Theorem 1.3 compares, in particular, positive supermodular dependent random vectors and also covers the extreme cases of positive supermodular dependence, i.e., independence and comonotonicity: Exactly in the case where $X_{i}$ and $X_{j}$ are independent for all $e=(i,j)\in E\,,$ we have that $(X_{n})_{n\in N}$ is a sequence of independent random variables. In the case of continuous marginal distributions, if $(X_{i},X_{j})$ is comonotonic for all $e=(i,j)\in E\,,$ then also $(X_{n})_{n\in N}$ is comonotonic. Note that comonotonicity models perfect positive dependence and relates to the upper Fréchet bound which is the greatest element with respect to the supermodular order in Frèchet classes (i.e., in classes of distributions with fixed marginals), see (12). For discontinuous marginal distributions, comonotonicity can generally not be modeled by a Markov tree distribution because the marginal distributions can affect the dependence structure in Markovian models, see Example A.2.
@@
    Theorem 1.3 extends comparison results for conditionally independent factor models and for discrete time Markov processes to Markov tree distributions, compare Figure 2. To see this, let $T$ be a chain, i.e., the edges of the tree are given by $E=\{(0,1),(1,2),(2,3),...\}\,.$ Then condition (i) in Theorem 1.3 is $X_{i+1}\uparrow_{st}X_{i}$ for $i\in\{1,2,3,\ldots\}$ and condition (ii) simplifies to $Y_{i}\uparrow_{st}Y_{i+1}$ for $i\in\{0,1,2,\ldots\}\,.$ Hence, Theorem 1.3 generalizes Proposition 1.1. In particular, we obtain that condition (i) in Proposition 1.1 can be skipped for $i=0\,.$ In the case where $T$ is a star on $d+1$ nodes, i.e., when $L=N\setminus\{0\}$ (all nodes except the root are leaves), then the set of edges is given by $E=\{(0,1),(0,2),\ldots,(0,d)\}\,.$ In this case, conditions (i) and (ii) of Theorem 1.3 simplify to $X_{j}\uparrow_{st}X_{0}$ for $i\in\{1,\ldots,d\}\setminus\{k^{*}\}$ and $Y_{j}\uparrow_{st}Y_{0}$ for $j\in\{1,\ldots,d\}\setminus\{\ell\}\,.$ Hence, Theorem 1.3 also generalizes Lemma 1.2 noting that condition (i) can be skipped for $i=k^{*}$ and condition (ii) can be skipped for $i=\ell\,.$
1.2 Related literature
The significance of Theorem 1.3 lies in the fact that it compares positive dependent multivariate distributions with respect to the strong notion of supermodular order. In the context of stochastic processes and time series, random variables in temporal or spatial proximity typically depend positively on each other. In risk management, for example, loan defaults are often positively dependent, or insurance losses exhibit positive dependencies [79]. In order to model positive dependence structures, comparison results are of particular importance, indicating the strength of the positive interrelations. Comparison results and various concepts of positive dependence are studied in the literature on multivariate parametric models for the normal distribution [82, 96, 106], for elliptical distributions [1, 7, 61, 120], for mixtures of elliptical distributions [90, 91], and for Archimedean copula models [86] (which corresponds to $l^{1}$-norm symmetric distributions [78]). For non-parametric distributions, general inequalities for positive dependent random variables are given under some structural assumptions, such as conditional independence, common marginals or exchangeability in [6, 8, 21, 105, 111, 112, 114]. Comparison results for Markov chains and Markov processes which exhibit positive dependencies are studied in [14, 16, 38, 57, 58, 69, 103]. The present paper contributes to the literature by extending various ordering results for discrete time Markov processes to tree structures and thus by providing general inequalities for Markov tree distributions. In particular, we obtain distributional robustness for various functionals, which we illustrate in the special case of hidden Markov models. For an overview of inequalities for multivariate distributions, see, e.g., [76, 87, 104, 113].
From a practical point of view, the supermodular order is of great importance in financial and actuarial mathematics. For instance, numerous payoff functions of financial derivatives are supermodular, see Table 1 and [9, 73, 110]. Further, by (10), the supermodular order is useful in risk management in quantifying portfolio risk and determining portfolio risk bounds under dependence information [87, 102], where the marginals can often be inferred from data; see, e.g., [23, 27, 109] and the references therein. From a theoretical point of view, inequalities for distributions with fixed marginals are studied in the field of optimal transport [92, 115]. In this regard, Theorem 1.3 yields solutions to optimal transport problems with additional structural information, such as the Markov property and positive dependencies. The supermodular order also allows for a clear interpretation because it can be described by simple rearrangements or mass transfers, as studied in [83].
In applications, typically, the one-dimensional, and to some extent, the two-dimensional marginal distributions can be estimated or partially estimated from data. However, due to the curse of dimensionality, the entire dependence structure of a random vector can generally not be determined. Therefore, models are needed that are flexible and robust on the one hand, but also easy to understand on the other hand. A large class of time series and evolution models are Markovian, where future and past events are independent conditionally on the present. Such models are completely specified by the bivariate distributions of adjacent variables, i.e., by the univariate marginal distributions and the bivariate copulas specifying the edges of a chain of variables, cf. Figure 1a). Nested models [62] allow for incorporating higher-order conditional dependencies, see [56, 77] for nested Archimedean copula models and [17, 36, 37] for vine copula models. The latter class of models allows to incorporate dependencies in a flexible way and is used in various applications, for example, in the fields of climate and wind [32, 46, 52, 60, 116], finance and risk management [59, 117, 119], and statistical learning [24, 72, 108]. However, further research on distributional and statistical properties of vine copula models is needed, see, e.g., [2, 40, 51, 81]. Since absolutely continuous Markov tree distributions are vine copula models truncated after the first level [26], the results presented in this paper also provide new insight into distributional properties of regular vine copula models.
@@ 1.2 Related literature
The rest of the paper is organized as follows. Section 2 provides the definitions of Markov tree distributions, the stochastic orderings, and positive dependence concepts, which we use in this paper. As a consequence of our main result, we derive in Section 3 various comparison results for the lower orthant, upper orthant, and directionally convex order, which also allow flexibility in the marginal distributions. Section 4 provides distributional robustness of various functionals on classes of hidden Markov models. A detailed discussion of the assumptions of our main result and a special ordering property for star structures in consistency with Chatterjee’s rank correlation are provided in Section 5. All proofs and, in particular, the proof of Theorem 1.3, which requires further technical details, are deferred to the appendix.
2 Preliminaries
@@
This section provides the basic notation and concepts used in this paper. It covers the definition of trees and copulas, which serve as the basic elements for constructing Markov tree distributions. Proposition 2.5 provides a simple representation of Markov tree distributions in terms of bivariate tree specifications, which is, on the one hand, a useful tool for constructing Markov tree distributions and, on the other hand, helpful to formalize the proofs of the comparison results studied in this paper. The second and third part of this section outline the definitions and basic relationships of the relevant stochastic orders and concepts of positive dependence. For the proofs of the results in this section, we refer to Appendix C.
@@ 2.1 Markov tree distributions
Trees can be used to model simple dependencies between random variables. While each node of a tree represents a random variable, the edges model the dependence structure between adjacent random variables [37, 68]. Markov tree distributions are uniquely determined by specifying univariate distribution functions associated with the nodes and bivariate copulas associated with the edges of the tree.
2.1.1 Trees and Markov realizations
We denote by $N$ an at most countable set of nodes which we label with the integers $N=\{0,1,\dots,d\}$ for $d\in\mathbb{N}\,,$ whenever $N$ has finitely many elements, and with $N=\mathbb{N}_{0}:=\{0\}\cup\mathbb{N}$ otherwise. We assume $|N|\geq 2$ to avoid cumbersome notation for trivial cases, where $|N|$ denotes the number of elements of $N\,.$ A graph on $N$ is a tuple $(N,E),$ where $E\subset N\times N$ is a set of oriented edges. By abuse of notation we write $\{i,j\}\in E$ if $(i,j)\in E$ or $(j,i)\in E$.
Definition 2.1 (Directed path, undirected path)
Let $(N,E)$ be a graph and let $i,j\in N$, $i\neq j$, be two nodes.
  1. (i)
    A directed path from $i$ to $j$ is a vector $(i,i_{1},\dots,i_{m},j)\subseteq N^{m+2}$ of $m+2$ nodes, $m\in\mathbb{N}_{0}\,,$ such that $(i,i_{1}),(i_{1},i_{2}),\dots,(i_{m},j)\in E$,
  2. (ii)
    A (undirected) path between $i$ and $j$ is a set $\{i_{1},\dots,i_{m}\}\subseteq N$ of $m$ distinct nodes, $m\in\mathbb{N}_{0}\,,$ such that $\{i,i_{1}\},\{i_{1},i_{2}\},\dots,\{i_{m},j\}\in E$,
In the literature on dependence modeling, trees are often defined as acyclic graphs, where the edges are unordered pairs of nodes [17, 35, 36, 63]. Since we generally allow asymmetric dependence properties, we focus on directed trees (a.k.a. polytrees or oriented trees) with a root that we label without loss of generality as $0\in N\,.$ Due the following definition, a tree is a graph in which all nodes can be reached from the root by a unique directed path. Such trees are also called arborescences.
Definition 2.2 (Tree)
A directed tree is a graph $T=(N,E)$ such that
  1. (i)
    for all $i\in N\backslash\{0\}$ there exists a unique directed path from 00 to $i$,
  2. (ii)
    $(i,j)\in E$ implies $(j,i)\notin E$,
In the following we refer to trees in the context of directed trees in the sense of Definition 2.2. By the definition of a tree, a node may have infinite degree, i.e., an infinite number of adjacent nodes. Due to (i) and (ii) in the above definition, an undirected path between $i$ and $j$ is uniquely determined and may contain the root. We denote this path by $p(i,j)\subseteq N\,$ (or equivalently by $p(j,i)\subseteq N$). By the definition of an undirected path, the nodes $i$ and $j$ are not included in $p(i,j)$. The leaves of a tree are defined as the subset of nodes in $N\backslash\{0\}$ having only one adjacent node, i.e.,
@@ 2.1.1 Trees and Markov realizations
[…]
Definition 2.3 (Markov tree dependence; [80, Definition 5])
Let $T=(N,E)$ be a tree. A distribution $\mu$ on $\mathbb{R}^{|N|}$ (resp. $\mathbb{R}^{\mathbb{N}_{0}}$ if $N=\mathbb{N}_{0}$) has Markov tree dependence (or is a Markov tree distribution) with respect to $T$ if there exists a sequence $(X_{i})_{i\in N}\sim\mu$ of random variables such that for every two finite disjoint subsets $A,B\subset N$ and for every $i\in N$ that separates $A$ and $B$, the vectors $X_{A}=(X_{j})_{j\in A}$ and $X_{B}=(X_{j})_{j\in B}$ are conditionally independent given $X_{i}$.
Weaker (i.e., non-Markovian) concepts of tree dependence can also be found in [17, 80]. Hierarchical tree structures which allow the modeling of higher order conditional dependencies are used in the context of vine copula models [17, 36, 37], noting that vine copula models truncated after the first level tree are Markov tree distributions [26].
2.1.2 Bivariate tree specifications
For various comparison results, we make use of the concept of copulas which is a tool that allows to study dependence structures between random variables. More precisely, a $d$-copula is a $d$-variate distribution function $C\colon[0,1]^{d}\to[0,1]$ having uniformly on $[0,1]$ distributed univariate marginal distributions. Due to Sklar’s theorem, every $d$-variate distribution function $F$ can be decomposed into its marginal distribution functions $F_{i}$, $1\leq i\leq d$, and a $d$-copula $C$ such that the joint distribution function can be expressed as the concatenation of these, i.e.,
@@ 2.1.2 Bivariate tree specifications
In this case $C$ is called a copula of $F$. The copula $C$ is uniquely determined on $\mathsf{Ran}(F_{1})\times\cdots\times\mathsf{Ran}(F_{d})\,,$ where $\mathsf{Ran}(f)$ denotes the range of a function $f\,.$ Further, for any copula $C$ and for any marginal distribution functions $F_{1},\ldots,F_{d}\,,$ the right-hand side of (5) defines a $d$-variate distribution function. If $X=(X_{1},\dots,X_{d})$ has distribution function $F\,,$ we say that $C=C_{X}$ is a copula of $X$. We denote by $\mathcal{C}_{d}$ the class of $d$-variate copulas. For an overview of the concept of copulas, see, e.g., [42, 88, 101].
As a consequence of the definition of Markov tree dependence, for any path $p(i,j)=\{i_{1},\ldots,i_{m}\}$ from $i$ to $j\,,$ the conditional distribution $X_{i}\mid(X_{i_{1}},\ldots,X_{i_{m}},X_{j})$ depends only on the random variable $X_{i_{1}}$, which is adjacent to $X_{i}\,.$This implies that Markov tree distributions are completely specified by bivariate distributions corresponding to the edges of the underlying tree. Due to Sklar’s theorem, each such bivariate distribution can be decomposed into two marginal distributions and a bivariate copula which describes the dependence structure. For a fixed tree, a bivariate tree specification assigns a univariate distribution function to each node and a bivariate copula to each edge of the tree as follows.
Definition 2.4 (Bivariate tree specification; [80, Definition 4])
@@
For a probability distribution $\mu$ on $\mathbb{R}^{|N|}\,,$ if $d<\infty\,,$ and on $\mathbb{R}^{\mathbb{N}_{0}}$, if $N=\mathbb{N}_{0}$, denote by $\mu_{n}$ and $\mu_{ij}\,,$ $n,i,j\in N\,,$ the univariate and bivariate marginal distributions with respect to the components $n$ and $(i,j)\,,$ respectively. Then $\mu$ is said to realize a bivariate tree specification $\mathcal{T}=(F,T,B)\,,$ if for all $n\in N$ and $e=(i,j)\in E\,,$ $F_{n}$ is the distribution function of $\mu_{n}$ and $B_{e}$ is a copula of $\mu_{ij}\,.$
Due to the following proposition, for every bivariate tree specification there exists a unique realizing Markov tree distribution, see [17] for the case when Lebesgue densities exist.
Proposition 2.5 (Markov realization of bivariate tree specification)
@@
We denote the uniquely determined Markov realization of a bivariate tree specification $\mathcal{T}=(F,T,B)$ by $\mathcal{M}(F,T,B)$ or $\mathcal{M}(\mathcal{T})$ and write $X\sim\mathcal{M}(F,T,B)$ for a sequence $X=(X_{n})_{n\in N}$ of random variables with Markov tree dependence specified by $\mathcal{T}\,.$
In the case that the marginal distributions and the bivariate copulas of a bivariate tree specification admit Lebesgue densities, the corresponding Markov realization has also a Lebesgue-density with a simple representation as follows, see [17, 80].
Proposition 2.6 (Density representation of Markov tree distributions)
Let $T=(N,E)$ be a tree with $d:=|N|<\infty$ and let $\mathcal{T}=(F,T,B)$ be a bivariate tree specification. Assume that $F_{n}$ and $B_{e}$ have Lebesgue densities $f_{n}$ and $b_{e}\,,$ respectively, for all $n\in N$ and $e\in E\,.$ Then, the Markov tree distribution $\mathcal{M}(F,T,B)$ has a Lebesgue density $g\colon\mathbb{R}^{d+1}\to[0,\infty)$ which is given by
${}g(x)=\prod_{n\in N}f_{n}(x_{n})\,\!\!\!\prod_{e=(i,j)\in E}b_{e}(F_{i}(x_{i}),F_{j}(x_{j}))\quad\text{for all }x=(x_{0},\ldots,x_{d})\in\mathbb{R}^{d+1}\,.$ (6)
2.2 Stochastic orderings

### v3

@@ 1.1 Main results
    The SI assumption in Theorem 1.3 are illustrated in Figure 2. While $X$ is SI from the root towards the leaves, the SI conditions for $Y$ hold in both directions, except along the predefined path $P$, where they are directed towards the root, and except at the leaves, where the SI conditions apply outward towards the leaves. Further, for the edge that lies on the path $P$ and connects to the leaf, no SI condition is required for $Y.$ Similarly, for an edge that connects the root with one of its children that, if possible, does not lie on the path $P$, no SI condition is required for $X.$
@@
    Theorem 1.3 extends the comparison results for conditionally independent factor models and for discrete-time Markov processes to Markov tree distributions: If $T$ is a chain, i.e., the edges of the tree are given by $E=\{(0,1),(1,2),(2,3),$...\}, then condition (i) in Theorem 1.3 is $X_{i+1}\uparrow_{st}X_{i}$ for $i\in\{1,2,3,\ldots\}$ and condition (ii) simplifies to $Y_{i}\uparrow_{st}Y_{i+1}$ for $i\in\{0,1,2,\ldots\}.$ Hence, Theorem 1.3 generalizes Proposition 1.1. In particular, we obtain that condition (i) in Proposition 1.1 can be skipped for $i=0,$ when we forgo with PSMD of $X.$ In the case where $T$ is a star on $d+1$ nodes, i.e., when $L=N\setminus\{0\}$ (all nodes except the root are leaves), then the set of edges is given by $E=\{(0,1),(0,2),\ldots,(0,d)\}.$ In this case, conditions (i) and (ii) of Theorem 1.3 simplify to $X_{j}\uparrow_{st}X_{0}$ for $i\in\{1,\ldots,d\}\setminus\{k^{*}\}$ and $Y_{j}\uparrow_{st}Y_{0}$ for $j\in\{1,\ldots,d\}\setminus\{\ell\}.$ Hence, Theorem 1.3 also generalizes Lemma 1.2 noting that condition (i) can be omitted for $i=k^{*}$, and condition (ii) can be omitted for $i=\ell.$ In this case, $X$ and $Y$ are, in general, no longer PSMD.
  5. (e)
    Corollary 1.4 provides a simple method to construct and compare PSMD Markov tree distributions: For any set of univariate distribution functions assigned to the nodes and for any set of bivariate SI copulas assigned to the edges of the tree, the implied Markov tree distribution is PSMD. Further, Corollary 1.4 yields a flexible method to construct multivariate parametric distributions that are increasing in all of their parameters with respect to the supermodular order. For fixed $F$ and $T$, this construction relies only on families of pointwise increasing CI copulas assigned to the edges of the tree $T$. While existing construction methods from the literature rely on the fact that the supermodular order is closed under mixtures or under independent or comonotonic concatenations [69, Theorem 9.A.3], our method is based on conditional independence along trees.
1.2 Structure of the paper
@@
The rest of the paper is organized as follows. Section 2 introduces Markov tree distributions and presents the relevant dependence concepts used throughout the paper. As a consequence of our main result, Section 3 provides comparison results when the marginals are in stochastic or convex order. Section 4 discusses, for star-shaped structures, a considerable extension of our supermodular ordering results to non-positive dependence specifications. Section 5 examines distributional robustness of hidden Markov models with a particular focus on the maximum of a perturbed random walk under model uncertainty. Various counterexamples that demonstrate the generality of our results are presented in Section 6. A detailed discussion of the assumptions of Theorem 1.3 is given in Appendix A. All proofs, including the technical proof of Theorem 1.3, are deferred to Appendix B.
2 Notation and basic dependence concepts
@@
This section provides the notation and dependence concepts used throughout the paper. Proposition 2.4 gives a simple representation of Markov tree distributions in terms of bivariate tree specifications. This representation serves as both a practical tool for constructing Markov tree distributions and a formal framework for various comparison results studied in this paper. The second part of this section covers the definitions and basic relations of the relevant stochastic orders and positive dependence concepts.
@@ 2.1 Markov tree distributions
Trees can be used to model simple dependencies between random variables. Each node of a tree represents a random variable, while the edges capture the dependence structure between random variables that are adjacent in the tree [23, 50, 49]. As we outline in the sequel, under the Markov property for tree structures, there exists a unique distribution on $\mathbb{R}^{N}$—the Markov tree distribution—fully determined by the univariate distribution functions assigned to the nodes and by the bivariate copulas assigned to the edges of the tree.
2.1.1 Trees and Markov tree dependence
We denote by $N$ an at most countable set of nodes which we label with the integers $N=\{0,1,\dots,d\}$ for some $d\in\mathbb{N}=\{1,2,\ldots\},$ whenever $N$ has finitely many elements, and with $N=\mathbb{N}_{0}:=\{0\}\cup\mathbb{N}$ otherwise. We assume $|N|\geq 2$ to avoid cumbersome notation for trivial cases, where $|N|$ denotes the number of elements of $N.$ A graph on $N$ is a tuple $(N,E),$ where $E\subset N\times N$ is a set of oriented edges. By abuse of notation we write $\{i,j\}\in E$ if $(i,j)\in E$ or $(j,i)\in E$. For $i,j\in N$ with $i\neq j,$ a directed path from $i$ to $j$ is a vector $(i,i_{1},\dots,i_{m},j)\subseteq N^{m+2}$ of $m+2$ distinct nodes, $m\in\mathbb{N}_{0},$ such that $(i,i_{1}),(i_{1},i_{2}),\dots,(i_{m},j)\in E$. Similarly, a (undirected) path between $i$ and $j$ is a set $\{i_{1},\dots,i_{m}\}\subseteq N$ of $m$ distinct nodes, $m\in\mathbb{N}_{0},$ such that $\{i,i_{1}\},\{i_{1},i_{2}\},\dots,\{i_{m},j\}\in E$.
In the literature on dependence modeling, trees are sometimes defined as acyclic graphs, where the edges are unordered pairs of nodes [16, 20, 23, 43]. However, since we generally allow for asymmetric dependence properties, we focus on directed trees (a.k.a. oriented trees or arborescences) with a designated root, which we label, without loss of generality, as $0\in N.$ According to the following definition, a tree is a graph in which all nodes can be reached from the root via a unique directed path.
Definition 2.1 (Directed tree)
A (directed) tree is a graph $T=(N,E)$ such that for all nodes $i\in N\backslash\{0\}$ there exists a unique directed path from 0 to $i$.
[…]
@@ 2.1.1 Trees and Markov tree dependence
The concept of Markov tree dependence uses a tree to model conditional independence between random variables indexed by the nodes of the tree; see [20, 50, 54]. Special cases are Markov processes in discrete time and conditionally independent factor models, where the underlying tree is a chain and a star, respectively (see Figure 1 a and b). A node $i\in N$ is said to separate two disjoint sets $A,B\subset N$ if for every $a\in A$ and $b\in B$ the path between $a$ and $b$ contains $i$.
Definition 2.2 (Markov tree dependence; [54, Definition 5])
Let $T=(N,E)$ be a tree. A distribution $\mu$ on $\mathbb{R}^{|N|}$ (resp. $\mathbb{R}^{\mathbb{N}_{0}}$ if $N=\mathbb{N}_{0}$) has Markov tree dependence (or is a Markov tree distribution) with respect to $T$ if there exists a sequence $(X_{n})_{n\in N}\sim\mu$ of random variables such that for every two finite disjoint subsets $A,B\subset N$ and for every $i\in N$ that separates $A$ and $B$, the vectors $X_{A}=(X_{j})_{j\in A}$ and $X_{B}=(X_{j})_{j\in B}$ are conditionally independent given $X_{i}$.
Weaker (i.e., non-Markovian) and stronger (i.e., higher-order Markovian) concepts of tree dependence are also discussed in [16, 54].
2.1.2 Bivariate tree specifications and Markov realizations
For various comparison results, we use the concept of copulas, a tool that allows us us to study dependence structures between random variables in full generality. More precisely, a $d$-copula is a $d$-variate distribution function $C\colon[0,1]^{d}\to[0,1]$ having uniformly on $[0,1]$ distributed univariate marginal distributions. Due to Sklar’s theorem, every $d$-variate distribution function $F$ on $\mathbb{R}^{d}$ can be decomposed into its univariate marginal distribution functions $F_{i}$, $1\leq i\leq d$, and a $d$-copula $C$ such that the joint distribution function can be expressed as the concatenation of these, i.e.,
@@ 2.1.2 Bivariate tree specifications and Markov realizations
In this case, $C$ is called a copula of $F$. The copula $C$ is uniquely determined on $\mathsf{Ran}(F_{1})\times\cdots\times\mathsf{Ran}(F_{d}),$ where $\mathsf{Ran}(f)$ denotes the range of a function $f.$ Further, for any copula $C$ and for all univariate distribution functions $F_{1},\ldots,F_{d},$ the right-hand side of (3) defines a $d$-variate distribution function. If $X=(X_{1},\dots,X_{d})$ has distribution function $F,$ we say that $C=C_{X}$ is a copula of $X$. For an overview of the concept of copulas, see e.g. [28, 62, 67].
As a consequence of the definition of Markov tree dependence (Definition 2.2), for any path $p(i,j)=\{i_{1},\ldots,i_{m}\}$ from $i$ to $j,$ the conditional distribution $X_{i}\mid(X_{i_{1}},\ldots,X_{i_{m}},X_{j})$ depends only on the random variable $X_{i_{1}}$, which is adjacent to $X_{i}.$This implies that Markov tree distributions are completely specified by the distributions of the bivariate random variables $(X_{k},X_{\ell}),$ $(k,l)\in E,$ that are adjacent in the underlying tree $T.$ Due to Sklar’s theorem, each such bivariate distribution function $F_{k\ell}=F_{X_{k},X_{\ell}}$ can be decomposed into its marginal distribution functions $F_{k}$ and $F_{\ell}$ and a bivariate copula $C_{k\ell}$, which describes the dependence structure between $X_{k}$ and $X_{\ell}$, i.e., $F_{k\ell}(x,y)=C_{k\ell}(F_{k}(x),F_{\ell}(y))$ for all $(x,y)\in\mathbb{R}^{2}.$ For a fixed tree, a bivariate tree specification assigns a univariate distribution function to each node and a bivariate copula to each edge of the tree as follows.
Definition 2.3 (Bivariate tree specification; [54, Definition 4])
@@
For a probability distribution $\mu$ on $\mathbb{R}^{|N|},$ if $d<\infty,$ or on $\mathbb{R}^{\mathbb{N}_{0}}$, if $N=\mathbb{N}_{0}$, denote by $\mu_{n}$ and $\mu_{ij},$ $n,i,j\in N,$ the univariate and bivariate marginal distributions with respect to the components $n$ and $(i,j),$ respectively. Then $\mu$ is said to realize a bivariate tree specification $\mathcal{T}=(F,T,B),$ if for all $n\in N$ and $e=(i,j)\in E,$ $F_{n}$ is the distribution function of $\mu_{n}$ and $B_{e}$ is a copula of $\mu_{ij}.$
Due to the following proposition, for every bivariate tree specification there exists a unique realizing Markov tree distribution; see [16] for the case when Lebesgue densities exist.
Proposition 2.4 (Markov realization of a bivariate tree specification)
@@
Recall that, for a bivariate tree specification $\mathcal{T}=(F,T,B),$ we denote its uniquely determined Markov realization by $\mathcal{M}(F,T,B)$, and we write $X\sim\mathcal{M}(F,T,B)$ for a sequence $X=(X_{n})_{n\in N}$ of random variables with this specification.
2.2 Stochastic orderings and positive dependence concepts